# State Your Intention to Steer Your Attention: An AI Assistant for Intentional Digital Living

### v2

@@ Abstract.
$c$s: Human-centered computing Empirical studies in HCIccs: Human-centered computing Interaction design
@@ 1. Introduction
Digital devices such as computers and smartphones have become essential tools for a wide range of tasks, spanning activities like studying, professional work, and planning travel. However, while using such devices, users are constantly assailed with tempting opportunities for distraction (Lyngs et al., 2019), both from the device itself and from their own thoughts of alternative on-device activities they could engage in. A variety of applications have been developed to help users stay on task (Roffarello and De Russis, 2023; Hiniker et al., 2016; Ko et al., 2015a; Kim et al., 2019; Collins et al., 2014; Ahmed et al., 2014). Yet, most remain limited to rule-based blockers or simple time trackers that operate within narrow boundaries. Such approaches fail to capture the context of what users intend to do versus what they are actually doing, and often impose excessive restrictions when the same application serves multiple purposes.
@@ 2. Related Work
Below we review three areas of related work: digital self-control tools, AI for digital well-being, and proactive assistants.
@@ 2.1. Digital Self-Control Tools
Digital self-control tools (DSCT) attempt to help people align their technology use with personal goals (Ko et al., 2015b; Roffarello and De Russis, 2023; Collins et al., 2014; Ahmed et al., 2014). Many prior studies have approached building DSCT with rule-based programs, such as allowing individuals or families to monitor usage against self-set limits (Hiniker et al., 2016; Ko et al., 2015a) or locking a device once a limit is exceeded (Kim et al., 2019). However, the rule-based DSCTs exhibit limitations, such as requiring manual configuration and a lack of adaptive interventions (Roffarello and De Russis, 2023). For example, a user might block YouTube with such a tool (because the website is typically a source of distraction) but then find the tool too inflexible to allow usage of YouTube for work tasks like learning about HCI (in contrast to the flexibility of INA illustrated in Figure 2). To address this, several studies attempt greater flexibility, such as varying the intensity or timing of interventions based on sensed mobile data like app usage history (Kovacs et al., 2018), allowing the relaxation of self-set rules like app blocks or time limits (Kovacs et al., 2021), and encouraging users to pause and reflect before engaging in potentially undesired actions (Cox et al., 2016). In this work, we propose a novel approach to building effective DSCTs: employing the remarkable capabilities of modern AI to detect the critical timing when a user’s activity diverges from their original intention and then providing timely, context-aware reminders.
2.2. AI for Digital Focus
Recently, various efforts have been made to leverage AI to reduce distraction. One of the attempts is building a prediction model of users’ off-task behavior while using digital devices from multi-modal data input, including app usage and screenshot (Chen et al., 2023). Another study reduced excessive smartphone use by locking the smartphone screen when overuse is detected. To unlock the screen, users are required to type specific digits, prompting a moment of reflection before continuing app usage. (Orzikulova et al., 2024) Other recent systems interact with users and improve their focus by utilizing large language models (LLMs) or conversational agents. For example, MindShift (Wu et al., 2024) intervenes when users attempt to open a blacklisted app by generating persuasive messages that discourage such usage. Also, StayFocused (Li et al., 2024) presents a specialized supportive application that blocks phone usage during a focus session, while providing reflective questions when the user tries to use it. While these approaches demonstrate the potential power of LLMs, their use of LLM is limited to message generation, and their detection of off-task behavior relies upon brittle rules. In contrast, our work aims to build a system that is context-aware in both its interactive messages and its off-task detection function, creating a deeply personalized experience.
2.3. Proactive Assistants
Proactive assistants are those that act without explicit user requests, shifting interactions from user-invoked to system-initiated while helping users follow through on their goals. Across many agent platforms, including robots (Darvish et al., 2025), mobile devices (Rawles et al., 2025; Lee et al., 2024b), and wearable devices (Fang et al., 2025), proactive assistants have been developed to provide various benefits to users, such as delivering timely, context-aware interventions and nudging users toward positive behavior change. For example, the Mirai system (Fang et al., 2025) provides contextually informed nudging, using a wearable camera and voice analysis to help users achieve pre-set goals such as “having a healthy diet” or “focusing on work.” However, the study primarily demonstrated the system’s technical feasibility through scenarios, rather than verifying its effectiveness on behavior change among real users with diverse goals. A recent study (Chen et al., 2025) introduced an LLM coding assistant that proactively provides suggestions to users when predefined, rule-based conditions are met. Building upon the tradition of these prior studies, we introduce a novel proactive assistant to promote intentional digital activities in general on-screen tasks.
3. Formative Study
We conducted formative interviews with individuals who use personal computers for long periods, with two goals: (1) to examine everyday digital activity and identify gaps between users’ intentions and their actual behavior, and (2) to explore the desirability of intelligent, context-aware interventions for supporting intentional digital activity. We describe the study procedure (Section 3.1), present key findings (Section 3.2), and outline the resulting design goals (Section 3.3).
@@ 3.1. Participants and Procedure
We recruited 11 participants (3 undergraduates, 4 graduate students, 4 office workers; ages 20–30, average age 25.6; 7 male, 4 female), whom we refer to as F1-F11. All participants reported using personal computers for at least four hours daily, providing a solid basis for examining discrepancies between intended and actual computer use. Eight participants had prior experience with productivity tools such as blocking or time-tracking applications, offering insights into the limitations of widely used conventional systems.
We conducted interviews focusing on three areas. First, we examined participants’ digital intentions: the goals they set for computer use, the strategies they employed to pursue these goals, and the patterns of deviation that emerged. Second, we investigated their experiences with existing productivity tools, including perceived benefits, limitations, and reasons for discontinuation. Third, we asked participants to envision an intelligent, context-aware assistant that could better support intentional digital activity, focusing on desired features, intervention styles, and timing. All interviews were audio-recorded with prior consent, manually transcribed, and analyzed using thematic coding. Participation was voluntary and uncompensated.
@@ 3.2. Findings
Participants described what an envisioned assistant for supporting intentional activity might look like. They preferred an assistant that could understand the user’s intention context, adapt flexibly to different situations, and intervene promptly when unintentional deviation from the stated intention was detected (F1, F9, F11). Rather than rigid blocking mechanisms, which were often perceived as intrusive and compulsory, participants expressed a strong preference for notification-based interventions that respected user autonomy. They also emphasized the importance of adjustable controls, such as fine-tuning how frequently the system should intervene or how persistent notifications should be (F3, F5, F6).
@@ 3.3. Design Goals
  • DG1: Intention Understanding. The system must first enable users to articulate their intentions. Because such intentions are often abstract and open to multiple interpretations, the system should clarify them through interaction. Clearer intentions allow the system to anticipate on-task actions more effectively and minimize interventions that are based on incorrect detection of distraction.
  • DG2: Context-Aware Distraction Detection. The system should understand the user activity flexibly, going beyond simple blacklist-based rules. To this end, it must intelligently analyze contextual cues (such as on-screen content) in relation to user intention to understand their activity continuously and adaptively.
  • DG3: Timely and Gentle Interventions. When a distraction is detected, the system should intervene promptly to help guide the user back to their original intention. Interventions should be polite and dismissible notifications rather than enforced restrictions. The system may also reinforce positive behaviors by offering praise or encouragement when the user returns to their intended activity.
  • DG4: Feedback-Driven Refinement.
    Given that any contemporary system will detect distractionimperfectly, the user should be able to provide feedback on its detection to improve its future detection accuracy. This feedback can be used not only to improve general detection but also user-specific detection, for cases when the correct judgment of distraction in one context depends on the specific user.
Figure 2. Overview of INA. (1) The process begins with a user’s initial, often abstract, intention. (2) The system engages in an LLM Q&A interaction to (3) establish a clarified user intention. (4) An LLM detector then analyzes on-screen activity, using a distraction score to classify the user’s state as on-task or off-task. (5) Based on this state, the system delivers a gentle nudge or positive reinforcement. (6) The user can provide feedback on this intervention, which (7) the LLM uses to refine its model, improving detection accuracy for future interactions. :
@@ 4.3. Supporting Users According to their Distraction State
Figure 3. (1-2) The user inputs an intention and presses ‘Set’ to initiate a session. (3-5) A chat interface is used for a brief Q&A clarification process, where the user responds to the LLM’s questions. (6) The user can also press ‘Start’ to skip this clarification and begin the session immediately. (7) During the session, INA displays the user’s current state with colored notifications: green for on-task and red for off-task. (8-9) The user can hover over a notification to provide feedback, marking it as correct or incorrect and optionally adding a reason. (10) The ‘Stop’ button can be pressed at any time to end the session. :
@@ 4.5. System Implementation
The implementation of INA was developed in Python. The client is a native macOS application built with PyQt to provide a standardized experimental environment, and the backend is implemented with FastAPI and deployed on Google Cloud Platform. LLM queries are processed through Gemini 2.0-Flash with a temperature of 0.1, while fixing other parameters to the default. To protect sensitive data, screenshots are masked using Cloud DLP (Data Loss Prevention) before upload, with images stored in Cloud Storage and corresponding metadata and event logs maintained in Firestore. Further technical details about implementation of INA are provided in Appendix A.5, and privacy and security considerations are discussed in Appendix E.
@@ 5.1. IntentionBench
Figure 4. An illustration of the IntentionBench construction process. We first collect focused sessions, where each session corresponds to executing a distinct task instruction (e.g., “Study HCI”). Then, two focused sessions are combined to synthesize a mixed session by concatenating and reordering their segments. The user intention is inherited from one session (on-task), while the other constitutes off-task segments, mimicking natural transition points of user workflows.:
@@ 5.2. Evaluation Results and Analysis
Figure 5. Overall procedure of our three-week, within-subjects field study. Participants attended a 30-minute online orientation session before the study. Participants used one application type on their computers each week. The assignment of each application type—INA, logging only, or simple reminder—was based on a random order. At the end of each week, participants completed a post-survey and a 10-20 minute semi-structured video interview.:
@@ 6.5. Methods of Analysis
Figure 6. (a) Mean LLM-estimated off-task ratio. (b) Mean intention alignment rating based on end-of-session self-reports. Error bars denote 95% confidence intervals across participants (user-averaged). ***, **, and *, indicate significance of p ¡ 0.001, p ¡ 0.01, p ¡ 0.05, respectively. Arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) values are more favorable.:
@@ 7.1. System Usage
Over the three-week study period, each participant used applications on average for 2.9 hours per day, amounting to roughly 20.6 hours per application across the three one-week conditions. This resulted in a total of 1,786 sessions, amounting to 1,360 hours of usage data and 2,449,690 system logs (screen images, app log information). On average, each session lasted 45.7 minutes, and sessions shorter than one minute were excluded from the analysis. Across simple reminder and INA sessions, participants entered 1,248 distinct intentions and provided 1,071 intention alignment ratings (85.8%) [^5] Survey data was not collected if the session ended prematurely due to application bugs or computer shutdowns. at the end of their sessions.
Furthermore, we examined participants’ feedback to INA. Out of 1,786 sessions, 19 participants provided explicit feedback during 90 sessions, resulting in 161 feedback instances. The majority of feedback was positive (111 ‘correct’, 68.9%), while 50 instances were negative (‘incorrect’, 31.1%). Notably, 42 feedback events contained free-form messages from users, allowing us to gain a deeper understanding of their experiences. Typical positive feedback messages expressed recognition and surprise (e.g., “accurate” and “you caught me”), while negative messages often explained the relevance of activities to their intentions or guided how the system should have worked.

### v3

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791404isbn: 979-8-4007-2278-3/2026/04submissionid: 1128ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Interaction design
@@ 1. Introduction
Digital devices such as computers and smartphones have become essential tools for a wide range of tasks, spanning activities like studying, professional work, and planning travel. However, while using such devices, users are constantly assailed with tempting opportunities for distraction (Lyngs et al., 2019), both from the device itself and from their own thoughts of alternative on-device activities they could engage in. A variety of applications have been developed to help users stay on task (Roffarello and De Russis, 2023; Hiniker et al., 2016; Ko et al., 2015; Kim et al., 2019; Collins et al., 2014; Ahmed et al., 2014). Yet, most remain limited to rule-based blockers or simple time trackers that operate within narrow boundaries. Such approaches fail to capture the context of what users intend to do versus what they are actually doing, and often impose excessive restrictions when the same application serves multiple purposes.
@@ 2. Related Work
In this section, we review relevant areas. These include the theoretical foundation of our work, existing digital self-control tools and their limitations, and recent AI-based systems in related domains.
2.1. Intention-Behavior Gap and Dual-Systems Theory
People frequently experience an intention–behavior gap (Sheeran and Webb, 2016), where their actions fail to align with their goals in domains such as health, academics, and work. This gap arises not only from limited willpower but from an interplay of factors, including insufficient execution plans, habitual behaviors, and environmental influences (Grimmer and Miles, 2016; Casais and Faria, 2021; Wee and Dillon, 2022). Repeated failures to bridge this gap can reduce self-efficacy and increase psychological stress (Adriaanse and ten Broeke, 2021; Chow et al., 2015). These challenges have intensified in the digital era, as engagement-driven platforms (Lyngs et al., 2020; Lukoff et al., 2021; Du et al., 2021) and constant notifications (Pielot and Rello, 2017; Okoshi et al., 2017) introduce persistent distractions that make it harder for people to follow through on their intentions in everyday digital life (Meier, 2021; Du et al., 2021).
We frame the intention–behavior gap in digital activities through Dual-Systems theory (Hofmann et al., 2009; Kahneman, 2003), which distinguishes fast, automatic reactions (System 1) from slow, deliberate control (System 2) (Stanovich and West, 2000). In digital environments, self-control failures arise when System 2 falters, allowing habitual System 1 responses (such as clicking notifications or recommendations) to override goal-directed intentions (Lyngs et al., 2019). Because System 2 is slow and resource-limited (Miller, 1956; Cowan, 2010), intention–behavior gaps readily emerge (Kotabe and Hofmann, 2015). In this work, we propose a method that provides timely intention reminders (Lyngs et al., 2019), thereby re-engaging System 2 to prevent automatic, unintentional System 1 behaviors from dominating.
@@ 2.2. Digital Self-Control Tools
Digital Self-Control Tools (DSCTs) assist users in reducing the intention–behavior gap in their digital activities (Roffarello and De Russis, 2023; Lyngs et al., 2019; Shen et al., 2019; Hiniker et al., 2016). A representative approach involves rule-based tools that block specific applications or websites or limit usage time (Hiniker et al., 2016; Ko et al., 2015). When users violate pre-set rules, these tools introduce friction, such as locking devices (Kim et al., 2019; Ko et al., 2016), displaying warning messages (Hiniker et al., 2016; Tseng et al., 2019), or requiring the typing of specific phrases (Cox et al., 2016), to discourage undesirable behavior. Another category of tools promotes self-reflection by visualizing usage patterns or displaying pop-ups that remind users of goals or tasks (Shen et al., 2019; Rooksby et al., 2016).
Despite their prevalence, most approaches are limited because they rely on static rules defined at the application or domain level (Roffarello and De Russis, 2023). Such static systems fail to distinguish between different digital contexts and therefore block access uniformly, often causing frustration (Lukoff et al., 2021; Oduor et al., 2016). Similarly, self-reflective pop-ups often appear as routine notifications unrelated to the user’s current context, leading users to ignore them and reducing their effectiveness (Kovacs et al., 2018, 2021). To address these limitations, recent work has explored leveraging Large Language Models (LLMs) to deliver more dynamic and personalized interventions (Wu et al., 2024; Li et al., 2024; Kim et al., 2024). However, these systems typically operate through chat-based interfaces, making it difficult to infer the user’s precise context or intervene seamlessly during ongoing tasks (Roffarello and De Russis, 2023; Okeke et al., 2018; Scibetta et al., 2025). In this work, we explore an AI-based, context-aware assistant that provides timely interventions based on users’ stated intentions and their on-screen activity.
2.3. AI-driven Intention Understanding
Recent advances in AI have demonstrated strong capabilities in understanding multimodal data and contextual information (Brown et al., 2020; Radford et al., 2021; Alayrac et al., 2022; Nguyen et al., 2025). In contrast to rule-based systems that depend heavily on costly and labor-intensive human specifications, AI-based approaches provide more adaptive and fine-grained mechanisms for interpreting user behaviors and inferring underlying intentions across diverse application domains, including context-driven LLM interfaces or situated AI systems (Jaber et al., 2024; Arakawa et al., 2025; Dang et al., 2025; Fang et al., 2025; Jones and others, 2024). For instance, Chen et al. (2023) developed a model capable of distinguishing time-killing behaviors on mobile devices with high precision. Shaikh et al. (2025) built a general user model for intention prediction, demonstrating a level of universality based on the common knowledge and context-awareness embedded in LLMs. These studies indicate that AI methods can function as reliable proxies for inferring user intent. Building on this insight, in this work, we leverage AI-based semantic alignment as an indicator of behavioral deviation and introduce a novel assistant system for digital self-control.
3. Formative Study for Assistant Design
While existing literature explains the phenomenon of digital distraction well, there is a lack of concrete guidelines on when and how an AI assistant should intervene within the screen-based digital activity context. To gain design insights for concrete functionalities and interaction strategies for such an assistant, we conducted formative interviews with individuals who use personal computers for their major activities. We describe the study procedure (Section 3.1), present key findings (Section 3.2), and outline the resulting design goals (Section 3.3).
@@ 3.1. Participants and Procedure
We recruited 11 participants (3 undergraduates, 4 graduate students, 4 office workers; ages 20–30, average age 25.6; 7 male, 4 female), whom we refer to as F1-F11. Participants were recruited individually through word-of-mouth within a university community. All participants reported using personal computers for at least four hours daily. We determined this group suitable for uncovering in-depth contexts and concrete needs for assistant design, considering its characteristic of being highly accustomed to digital environments. Eight participants had prior experience with productivity tools such as blocking or time-tracking applications, offering insights into the limitations of widely used conventional systems.
We conducted interviews focusing on three areas. First, we examined participants’ digital intentions: the goals they set for computer use, the strategies they employed to pursue these goals, and the patterns of deviation that emerged. Second, we investigated their experiences with existing productivity tools, including perceived benefits, limitations, and reasons for discontinuation. Third, we asked participants to envision an intelligent, context-aware assistant that could better support intentional digital activity, focusing on desired features, intervention styles, and timing. All interviews were audio-recorded with prior consent, manually transcribed. We analyzed the data using inductive thematic analysis (Braun and Clarke, 2006). Two researchers independently coded the transcripts and refined the themes through iterative discussion. Participation was voluntary and uncompensated.
@@ 3.2. Findings
Participants described what an envisioned assistant for supporting intentional activity might look like. They preferred an assistant that could understand the user’s intention context, adapt flexibly to different situations, and intervene promptly when unintentional deviation from the stated intention was detected (F1, F9, F11). Rather than rigid blocking mechanisms, which were often perceived as intrusive and compulsory, participants expressed a strong preference for notification-based interventions that respected user autonomy. They also emphasized the importance of adjustable controls, such as fine-tuning how frequently the assistant should intervene or how persistent notifications should be (F3, F5, F6).
@@ 3.3. Design Goals
  • DG1: Intention Understanding. The assistant must first enable users to articulate their intentions. Because such intentions are often abstract and open to multiple interpretations, the assistant should clarify them through interaction. Clearer intentions allow the assistant to anticipate on-task actions more effectively and minimize interventions that are based on incorrect detection of distraction.
  • DG2: Context-Aware Distraction Detection. The assistant should understand the user activity flexibly, going beyond simple blacklist-based rules. To this end, it must intelligently analyze contextual cues (such as on-screen content) in relation to user intention to understand their activity continuously and adaptively.
  • DG3: Timely and Gentle Interventions. When a distraction is detected, the assistant should intervene promptly to help guide the user back to their original intention. Interventions should be polite and dismissible notifications rather than enforced restrictions. The assistant may also reinforce positive behaviors by offering praise or encouragement when the user returns to their intended activity.
  • DG4: Feedback-Driven Refinement. Given that any contemporary assistant will detect distraction imperfectly, the user should be able to provide feedback on its detection to improve its future detection accuracy. This feedback can be used not only to improve general detection but also user-specific detection, for cases when the correct judgment of distraction in one context depends on the specific user.
Refer to caption A diagram showing the workflow of INA. When a user inputs an intention, INA elicits the intention, monitors the user behavior with a distraction detector module, provides supportive nudges, and improves accuracy through user feedback.
Figure 2. Overview of INA. (1) The process begins with a user’s initial, often abstract, intention. (2) The system engages in an LLM Q&A interaction to (3) establish a clarified user intention. (4) An LLM detector then analyzes on-screen activity every two seconds, using a distraction score to classify the user’s state as on-task or off-task. (5) Based on this state, the system delivers a gentle nudge or positive reinforcement. (6) The user can provide feedback on this intervention, which (7) the LLM uses to refine its model, improving detection accuracy for future interactions.
@@ 4.2. Monitoring User Behavior to Detect Distraction
Figure 3. (1-2) The user inputs an intention and presses ‘Set’ to initiate a session. (3-5) A chat interface is used for a brief Q&A clarification process, where the user responds to the LLM’s questions. (6) The user can also press ‘Start’ to skip this clarification and begin the session immediately. (7) During the session, INA displays the user’s current state with colored notifications: green for on-task and red for off-task. (8-9) The user can hover over a notification to provide feedback, marking it as correct or incorrect and optionally adding a reason. (10) The ‘Stop’ button can be pressed at any time to end the session.
Interface screenshot of INA, showing how a user sets an intention input, clarifies it through Q&A, receives on-task or off-task notifications, and provides feedback during the session.
@@ 4.4. Improving Accuracy Through User Interaction
Figure 4. An illustration of the IntentionBench construction process. We first collect focused sessions, where each session corresponds to executing a distinct task instruction (e.g., “Study HCI”). Then, two focused sessions are combined to synthesize a mixed session by concatenating and reordering their segments. The user intention is inherited from one session (on-task), while the other constitutes off-task segments, mimicking natural transition points of user workflows.
A diagram showing how IntentionBench is constructed by combining two focused sessions into a mixed session.
@@ 4.5. System Implementation
The implementation of INA was developed in Python. The client is a native macOS application built with PyQt to provide a standardized experimental environment, and the backend is implemented with FastAPI and deployed on Google Cloud Platform. LLM queries are processed through Gemini 2.0-Flash with a temperature of 0.1, while fixing other parameters to the default. To protect sensitive data, screenshots are masked using Cloud DLP (Data Loss Prevention) before upload, with images stored in Cloud Storage and corresponding metadata and event logs maintained in Firestore. Original screen images used for real-time inference are immediately discarded. Access to stored masked images and logs is restricted based on least-privilege IAM with multi-factor authentication, and all data are encrypted in transit and at rest. Further technical details about implementation of INA are provided in Appendix A.5, and privacy and security considerations are discussed in Appendix E.
@@ 6.2. Baseline Systems: Logging only and Simple Reminder Applications
Figure 5. Overall procedure of our three-week, within-subjects field study. Participants attended a 30-minute online orientation session before the study. Participants used one application type on their computers each week. The assignment of each application type—INA, logging only, or simple reminder—was based on a random order. At the end of each week, participants completed a post-survey and a 10-20 minute semi-structured video interview.
A diagram of the three-week field study, showing the procedure of orientation, weekly use of different applications, and post-survey with interviews.
@@ 6.5. Methods of Analysis
Figure 6. (a) Mean LLM-estimated off-task ratio. (b) Mean intention alignment rating based on end-of-session self-reports. Error bars denote 95% confidence intervals across participants (user-averaged). ***, **, and *, indicate significance of p ¡ 0.001, p ¡ 0.01, p ¡ 0.05, respectively. Arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) values are more favorable.
Bar charts comparing INA and simple reminder in terms of off-task ratio and intention alignment rating.
@@ 7.1. System Usage
Over the three-week study period, each participant used applications on average for 2.9 hours per day, amounting to roughly 20.6 hours per application across the three one-week conditions. This resulted in a total of 1,786 sessions, amounting to 1,360 hours of usage data and 2,449,690 system logs (screen images, app log information). On average, each session lasted 45.7 minutes, and sessions shorter than one minute were excluded from the analysis. Across simple reminder and INA sessions, participants entered 1,248 distinct intentions and provided 1,071 intention alignment ratings (85.8%) [^5] Survey data was not collected if the session ended prematurely due to application bugs or computer shutdowns. at the end of their sessions. Furthermore, we examined participants’ feedback to INA. Out of 1,786 sessions, 19 participants provided explicit feedback during 90 sessions, resulting in 161 feedback instances. The majority of feedback was positive (111 ‘correct’, 68.9%), while 50 instances were negative (‘incorrect’, 31.1%). Notably, 42 feedback events contained free-form messages from users, allowing us to gain a deeper understanding of their experiences. Typical positive feedback messages expressed recognition and surprise (e.g., “accurate” and “you caught me”), while negative messages often explained the relevance of activities to their intentions or guided how the system should have worked.